# Understanding and Improving Information Preservation in Prompt Compression for LLMs

### v1

@@ 5 Improving Soft Prompt Compression
In Section 4 we identify key weaknesses of soft prompt compression focusing on xRAG. We highlight its inability to preserve detailed information from the original input, especially in long-context scenarios. In this section, we explore directions to mitigate these limitations, while still retaining the advantage of high compression rates. We use our prompt compression evaluation framework to verify whether the introduced modifications lead to improvements over the original xRAG method. We re-train an xRAG model using the data and codebase released with its paper to ensure that the models are fully comparable (details are in Appendix D). Throughout this section, we denote our reproduced version as xRAG (Reproduced).
5.1 Sentence Pre-Training and Sentence-Level Fine-Tuning (xRAG w/ Sentence PT + FT)
@@
The only trainable component of the xRAG is a bridge model. Its main role is aligning two embedding spaces – the one from the encoder (dense retrieval) model, and the one from the target LLM. This function is learned via unsupervised pre-training on a reconstruction task, and then further fine-tuned on a set of target downstream tasks.
@@
In xRAG, the pre-training (PT) samples are encoded into single xRAG tokens (obtained from the bridge model), regardless of their length. Since the majority of the examples are longer than one sentence, [^5] Distribution of number of sentences per training sample in pre-training data can be found in Figure 2 in Appendix C.1. the bridge model fails to transform the representations of texts with smaller granularity (see Table 3), and thus fails to capture the details beyond the topic of the samples. To address this limitation, we use single-sentence samples in the pre-training to ensure that the model can effectively handle these basic cases. We use the same pre-training data, but only the first sentence from each sample is considered. In the fine-tuning (FT) phase, original xRAG chunks each context into texts of 180 tokens with each chunk encoded into one xRAG token. Instead, we replace token-based chunking with sentence-based segmentation with each sentence encoded into a separate xRAG token to preserve information continuity. We denote this variant of xRAG as xRAG w/ Sentence PT + FT.
@@
Figure 1: Information preservation results for different xRAG variants. Similarity is measured with BERTScore between the original and reconstructed text, while entity preservation is based on EM of entities in the reconstruction.
@@
On the downstream evaluation, we see that the xRAG w/ Sentence PT + FT brings sizable improvements on all tasks (see Table 6). With context level (Cont.) compression we see 1-5 points absolute gains on all tasks (up to 13% relative), suggesting that this method is able to capture the information better. This is further supported by the increase in the grounding of the generations (see Table 7). We also see that the model is now able to handle sentence-level compression better (Sent.) – the largest gains we see on HotpoQA* (downstream task performance increases from 0.174 to 0.423) and TriviaQA (from 0.115 to 0.361). Interestingly, the sentence-level representations are still not performing as well as the context-level ones.
@@
Finally, in the context reconstruction experiments, we see improved performance when encoding at sentence level. [^6] Reconstructed samples are in Table 23 in Appendix E.3. The xRAG w/ Sentence PT + FT) model is able to preserve more entities on average—increase from 13% to 20% (see Figure 1). However, in absolute terms, the numbers still remain low (preserving between 13%-39% entities). In all three dimensions, the performance on sentence-level tokens is lower than when encoding the entire context at once, implying that the model is not able to handle multiple tokens at a time.
5.2 Two-Step Pre-Training and Sentence-Level Fine-Tuning (xRAG w/ Two-Step PT + FT)
@@
Even though we see improvements in terms of downstream and grounding with the xRAG w/ Sentence PT + FT model (§ 5.1), the sentence-level encoding is still falling behind. We hypothesize that the main reason is that the model is not able to handle information spread across multiple tokens.
@@
To verify and refine this hypothesis, we introduce an additional pre-training step that focuses on helping model reason about information from multiple xRAG tokens. In particular, we adopt a two-stage pre-training procedure with the following steps: (i) encoding one sentence per pre-training sample, and (ii) chunking samples into sentences with each sentence encoded separately. We keep the fine-tuning stage unchanged as in xRAG w/ Sentence PT + FT, i.e., the samples are chunked into sentences with one token per sentence. We denote this variant as xRAG w/ Two-Step PT + FT.
@@
Results on downstream tasks show that prompt compression methods can further benefit from sentence-level encoding, particularly for datasets with focused, concise contexts like HotpotQA*, where low-level facts are needed for reasoning (see Table 6). We see 23% relative improvement over the baseline (from 0.375 EM for xRAG (Reproduced) Cont. to 0.462 EM for xRAG w/ Two-Step PT + FT Sent.). In general, we observe consistent improvements with this model compared to the baseline when the information is encoded on sentence level (Sent. columns). However, the model is still not able to fully process complex, noisy contexts, as evident in results for HotpotQA and TriviaQA, where the performance with one token per context drops (Cont. columns).
@@
Even though we observe lower grounding for multi-token encoded texts, the reconstruction experiments show that our two-step pre-training on data with different granularity results in significant improvements in terms of information preservation. Most importantly, improvements in the model’s capability to reconstruct compressed information are visible independently of the amount of encoding tokens used. On average, we can reconstruct 50% of entities from the original text encoded into multiple tokens (compared to 13% for the xRAG (Reproduced)) (see Figure 1). While cardinal entities remain the most challenging ones to preserve, the two-step model retains significantly more of them (0.36 vs. 0.23 at the context level and 0.50 vs. 0.09 at the sentence level). Moreover, we observe even higher similarity scores for text encoded into multiple tokens (F1 of 0.63 vs. 0.19), showing that our strategy captures finer-grained information.
@@ 5.3 Compression Rates
Table 8 presents the compression rates for each model (measured in words and including the system instructions, context, etc.). The sentence-level xRAG retains a high compression rate with a 4x-12x reduction in the input size and a 2x higher compression rate than LLMLingua. The compression rate depends on the lengths of the inputs but the compression granularity is a controllable parameter of the method.
@@
Table 8: Average prompt lengths (in words) after compression. The compression rate is shown in brackets.
@@ 6 Lessons Learned and Future Directions
So far, we have seen that compression techniques fail to preserve detailed information, especially in long-context scenarios and at high compression rates, leading to drops in performance and grounding. In line with this, context reconstruction experiments indicate that xRAG tokens primarily capture the general topic of the compressed content but fail to retain essential details for a task (numbers, dates, names, etc.). We propose modifications of xRAG and provide evidence that manipulating information granularity during training improves information retention in soft prompts, which translates to improved performance and grounding in downstream tasks. Notably, improvements occur regardless of the number of encoding tokens used.
@@
Given the complexity and information density of text in information-intensive tasks, an effective alternative to encoding everything into a single soft prompt token is needed (Mu et al., 2023; Li et al., 2024b). Our experiments suggest that while increasing text granularity is promising, reasoning across information stored independently in soft prompt tokens remains challenging. To address this, emphasizing multi-token sample training is crucial for improving the model’s ability to integrate information across separate units (Cheng et al., 2024). Enhancing the model’s capacity to process multi-token compression will also enable more advanced approaches, such as dynamic compression rates.
@@
Multi-token soft prompting is an important direction for compression methods. However, using a task-independent one-size-fits-all strategy does not capture the fact that every input comes with different complexity and level of details required for the target task Chevalier et al. (2023); Mu et al. (2023); Jiang et al. (2023b); Pan et al. (2024b). Moreover, despite the advancements in prompt compression methods, grounding and information retention remain suboptimal, especially when encoding with multiple tokens. Dynamically adjusting the compression rate and the compression strategy can improve contextual understanding on information spread in several tokens (Nagle et al., 2024).
@@
Preserving details, such as named entities, in compressed representations is crucial for improving performance on tasks that require reasoning. Hard compression Li et al. (2023); Jiang et al. (2023b); Chuang et al. (2024) is inherently explainable and can explicitly retain these details. In contrast, soft prompting has limited representational capacity, making it challenging to encode certain information, such as cardinals, in highly compressed embeddings Wallace et al. (2019); Jiang et al. (2020); Epure and Hennequin (2022). A promising future research direction is combining soft and hard prompting, where hard prompting preserves key details (e.g., numbers, names, etc.), while the remaining context goes into soft prompts, ensuring that the critical information is not lost during the compression.
@@ 7 Conclusions
In this paper, we propose a framework for assessing the quality of prompt compression methods on long-context generation tasks. We shed light on both the capabilities and limitations of two state-of-the-art hard (LLMLingual) and soft (xRAG) compression approaches. Our findings reveal that existing compression techniques fail to preserve detailed information, especially in long context scenarios. We explore several directions to improve the soft compression method, showing an increase in model performance – up to +23% on downstream tasks, up to 8 BERTScore points in grounding, and preserving 2.7x more entities.
@@
Our study has several limitations that should be considered when interpreting the results. First, all experiments are conducted using Mistral-7B as the target LLM. While this allows for consistent evaluation, exploring different model sizes and architectures is beyond the scope of this work and remains an area for future research. Second, our analysis focuses exclusively on English benchmarks. Findings may not generalize to multilingual settings, where language-specific characteristics could influence the effectiveness of prompt compression methods. Third, we evaluate only one representative baseline from each category of prompt compression techniques. While this provides an initial comparison, a broader investigation using our evaluation framework is needed to fully understand the trade-offs across different methods. Finally, although we examine multiple generation tasks, there is still room for further exploration, particularly in long-form QA scenarios, where prompt compression may have a more significant impact on model performance. Future analysis should also take into account the computational cost of achieving compression which is an important factor to be considered when comparing different methods.
@@ Appendix A Implementation Details
In all the experiments we use Mistral-7B Instruct v0.2 (Jiang et al., 2023a)777huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 that contains approximately 7.24B parameters as a target LLM. We use the associated model checkpoint implemented using the PyTorch (Paszke et al., 2019) framework that is available on the HuggingFace Hub via the transformers (Wolf et al., 2020) library. We access microsoft/phi-2 (Abdin et al., 2023) [^8] https://huggingface.co/microsoft/phi-2 for LLMLingua and the Mistral 7B based xRAG (Hannibal046/xrag-7b [^9] https://huggingface.co/Hannibal046/xrag-7b) in the same way. We load the datasets from the HuggingFace using the Datasets library Lhoest et al. (2021).
@@
We train our models using two compute instances, each equipped with 8 NVIDIA A100 Tensor Core GPUs, 1152 GiB of memory, and 96 vCPUs. The training process for a new xRAG model variant typically takes around 16 hours. We adopt a batch size of 1 per GPU and leverage the Accelerate library with a gradient accumulation step of 4.101010https://huggingface.co/docs/accelerate The specific hyper-parameters used during training are detailed in Table 9.
@@ C.1 Training Dataset
Figure 2: Distribution of the lengths of the examples measured in terms of number of sentences per training sample in pre-training data. The training partition contains 26,541,264 samples overall. The longest sample of which contains 137 sentences.
@@ C.2 Evaluation Datasets
We use five different datasets for downstream task evaluation: HotpotQA (Yang et al., 2018), arXiv-sumarization (Cohan et al., 2018), QuAC (Choi et al., 2018), TriviaQA (Joshi et al., 2017), and GSM8K (Cobbe et al., 2021). The distribution of context length for each dataset is presented in Figure 3.
@@ Appendix D Reproducibility of Baselines
For xRAG, we assess performance on 1,000 randomly selected samples from the HotpotQA development set (see Table 10). We evaluate two versions: (i) the vanilla xRAG model released by the authors,9 and (ii) our reproduced version trained using the original data and codebase (github.com/Hannibal046/xRAG). The reproduced model follows the same pre-training and fine-tuning setup, though we use approximately 50% of the fine-tuning samples with original context paragraphs instead of retrieved ones due to limitations in reproducing the original data pre-processing steps.

### v2

@@ 5 Improving xRAG
The results presented in Section 4 show a large disproportion between PISCO and the remaining two prompt compression methods. In particular, we highlight xRAG’s inability to preserve detailed information from the original input in long-context scenarios. We here explore directions to mitigate these limitations of xRAG, while still retaining the advantage of high compression rates. We use our prompt compression evaluation framework to verify whether the modifications lead to improvements over the original xRAG method. We re-train an xRAG model using the data and code released with its paper to ensure that the models are fully comparable (details in Appendix D). We denote this reproduced version as xRAG (Reproduced).
5.1 Sentence Pre-Training and Sentence-Level Fine-Tuning (w/ Sentence PT + FT)
@@
The only trainable component of the xRAG is a bridge model, whose main role is to align two embedding spaces – the one from the encoder (dense retrieval) model and the one from the target LLM. This function is learned via unsupervised pre-training on a reconstruction task, and then further fine-tuned on a set of target downstream tasks.
Figure 1: Distribution of the lengths of the examples measured in terms of number of sentences per training sample in pre-training data. The training partition contains 26,541,264 samples overall. The longest sample contains 137 sentences.
@@
In xRAG, the pre-training (PT) samples are encoded into single xRAG tokens, regardless of their length. Since the majority of the examples are longer than one sentence (see Figure 1), the bridge model fails to transform the representations of texts with smaller granularity (see Table 3) and thus to capture details beyond the topic of the samples. To address this limitation, we use single-sentence samples in pre-training to ensure that the model can effectively handle these basic cases. We use the same pre-training data, but consider only the first sentence from each sample. In the fine-tuning (FT) phase, original xRAG chunks each context into texts of 180 tokens with each chunk encoded into one xRAG token. Instead, we replace token-based chunking with sentence-based segmentation with each sentence encoded into a separate xRAG token to preserve information continuity. We denote this variant of xRAG as xRAG w/ Sentence PT + FT.
@@
Figure 2: Information preservation results for xRAG variants. Similarity is measured with BERTScore between the original and reconstructed text. Entity preservation is based on EM of entities in the reconstruction.
@@
On the downstream evaluation, xRAG w/ Sentence PT + FT brings sizable improvements on all tasks (Table 6). With context-level (Cont.) compression we see 1-5 points absolute gains on all tasks (up to 13% relative), suggesting that this method allows to capture the information better. This is further supported by the increase in grounding of the generated text (Table 9). We also see that the model is able to handle sentence-level compression better (Sent.) – especially on HotpotQA* (downstream performance increases from 0.174 to 0.423) and TriviaQA (from 0.115 to 0.361). Interestingly, sentence-level representations are still not performing as well as the context-level ones.
@@
Finally, in the context reconstruction experiments, we see improved performance when encoding at sentence level. [^4] Reconstructed samples are in Table 24 in Appendix E. The xRAG w/ Sentence PT + FT) model is able to preserve more entities on average—increase from 13% to 20% (see Figure 2). However, in absolute terms, the numbers still remain low (preserving between 13%-39% entities). In all three dimensions, the performance on sentence-level tokens is lower than when encoding the entire context at once, implying that the model is not able to handle multiple tokens at a time.
5.2 Two-Step Pre-Training and Sentence-Level Fine-Tuning (w/ Two-Step PT + FT)
@@
Even though we see improvements in downstream performance and grounding with the xRAG w/ Sentence PT + FT model (§ 5.1), the sentence-level encoding is still falling behind. We hypothesize that the main reason is that the model is not able to handle information spread across multiple tokens.
@@
To verify and refine this hypothesis, we introduce an additional pre-training step that focuses on helping the model reason about information from multiple xRAG tokens. We adopt a two-stage pre-training procedure with these steps: (i) encoding one sentence per pre-training sample; (ii) chunking samples into sentences with each sentence encoded separately. The fine-tuning stage is the same as in xRAG w/ Sentence PT + FT (chunking samples into sentences with one token per sentence). We denote this variant as xRAG w/ Two-Step PT + FT.
@@
Results on downstream tasks show that prompt compression methods can further benefit from sentence-level encoding, particularly for datasets with focused, concise contexts like HotpotQA*, where low-level facts are needed for reasoning (Table 6). We see 23% relative improvement over the baseline (in EM, from 0.38 for xRAG (Reproduced) Cont. to 0.47 for xRAG w/ Two-Step PT + FT Sent.). We generally observe consistent improvements over the baseline with this model when the information is encoded at sentence-level (Sent. columns). Still, the model cannot fully process complex, noisy contexts, as evident in results for HotpotQA and TriviaQA, where the performance with one token per context drops (Cont. columns).
@@
Even though we observe lower grounding for multi-token encoded texts, the reconstruction experiments show that our two-step pre-training on data with different granularity results in significant improvements in terms of information preservation. Crucially, this effect is visible independently of the amount of encoding tokens used. On average, we can reconstruct 50% of entities from the original text encoded into multiple tokens (compared to 13% for the xRAG (Reproduced)) (Figure 2). While cardinal entities remain the most challenging ones to preserve, the two-step model retains significantly more of them (0.36 vs. 0.23 at context-level and 0.50 vs. 0.09 at sentence-level). Moreover, we observe even higher similarity scores for text encoded into multiple tokens (F1 of 0.63 vs. 0.19), showing that our strategy captures finer-grained information.
5.3 Qualitative Analysis
To understand the type of information preserved, we conducted a qualitative analysis of reconstructed examples, examining 10 samples of varying lengths generated using 20 different prompts. Our investigation focused on prompts from the xRAG paraphrase training phase. The analysis confirmed our hypothesis that the xRAG tokens primarily capture the general topic of the compressed content but fail to retain specific details. This behavior is expected, as xRAG relies on a dense retrieval model trained for document similarity. Depending on the topic, the LLM can reconstruct the main entity described in the text, aligning with its strong performance on the TriviaQA dataset. However, numerical entities such as dates and numbers are often lost. Examples of reconstructed contexts can be found in Table 24 in Appendix E.
@@ 5.4 Compression Rates
Table 10 presents the compression rates for each model (measured in tokens and including the system instructions, context, etc.). PISCO achieves very high compression rates without performance loss in any of the aspects investigated with our evaluation framework, establishing it as the state-of-the-art for prompt compression in long-context scenarios. The sentence-level xRAG retains a high compression rate with a 4x-12x reduction in the input size and a 2x higher compression rate than LLMLingua. The compression rate depends on the lengths of the inputs but the compression granularity is a controllable parameter of the method.
@@
Table 10: Average prompt lengths (in tokens) after compression. The compression rate is shown in brackets. See Table 8 for compression computational efficiency.
@@ 6 Lessons Learned and Future Directions
So far, we have seen that many compression techniques fail to preserve detailed information, especially in long-context scenarios and at high compression rates, leading to drops in performance and grounding. In line with this, context reconstruction experiments indicate that xRAG tokens primarily capture the general topic of the compressed content but fail to retain essential details for a task (numbers, dates, names, etc.). We propose modifications of xRAG and provide evidence that manipulating information granularity during training improves information retention in soft prompts, which translates to improved performance and grounding in downstream tasks. Notably, improvements occur regardless of the number of encoding tokens used.
@@
Given the complexity and information density of text in information-intensive tasks, an effective alternative to encoding everything into a single soft prompt token is needed (Mu et al., 2023; Li et al., 2024b). Our experiments suggest that while increasing text granularity is promising, reasoning across information stored independently in soft prompt tokens remains challenging. PISCO addresses this by using sequence-level distillation, rather than token-level (Louis et al., 2025). Our results confirm that the training with multi-token samples implemented in PISCO significantly improves grounding and information preservation and is crucial for the ability to integrate information across separate units (Cheng et al., 2024).
@@
Multi-token soft prompting is an important direction for compression methods. However, a task-independent one-size-fits-all strategy does not capture the fact that every input comes with different complexity and level of detail required for the target task Chevalier et al. (2023); Mu et al. (2023); Jiang et al. (2023b); Pan et al. (2024b). Moreover, despite the advancements in prompt compression methods, grounding and information retention remain suboptimal, especially when encoding with multiple tokens. Dynamically adjusting the compression rate and the compression strategy can improve contextual understanding on information spread in several tokens (Nagle et al., 2024).
@@
Preserving details, such as named entities, in compressed representations is crucial for improving performance on tasks that require reasoning. Hard compression Li et al. (2023); Jiang et al. (2023b); Chuang et al. (2024) is inherently explainable and can explicitly retain these details. In contrast, soft prompting has limited representational capacity, making it challenging to encode certain information, such as cardinals, in highly compressed embeddings Wallace et al. (2019); Jiang et al. (2020); Epure and Hennequin (2022). A promising approach combines both: hard prompts preserve key details (e.g., numbers, names), while the rest of the context is encoded via soft prompts, reducing information loss during the compression.
@@ 7 Conclusions
In this paper, we propose a framework for assessing the quality of prompt compression methods on long-context generation tasks. We shed light on both the capabilities and limitations of the state-of-the-art hard (LLMLingua) and soft (xRAG and PISCO) compression approaches. Our findings reveal that some existing compression techniques fail to preserve detailed information, especially in long-context scenarios. We explore several directions to improve the xRAG soft compression method, showing an increase in model performance – up to +23% on downstream tasks, up to 8 BERTScore points in grounding, and preserving 2.7x more entities. [^5] The code repository: https://github.com/amazon-science/information-preservation-in-prompt-compression.
@@
Our study has several limitations that should be considered when interpreting the results. First, all experiments are conducted using Mistral-7B as the target LLM. While this allows for consistent evaluation, exploring different model sizes and architectures is beyond the scope of this work and remains an area for future research. Second, our analysis focuses exclusively on English benchmarks. Findings may not generalize to multilingual settings, where language-specific characteristics could influence the effectiveness of prompt compression methods. We believe that our findings will not change across languages, as we outline fundamental issues with the compression frameworks. If anything, we expect the soft approaches to be even more shallow in capturing details in other languages. Third, we evaluate only one representative baseline from each category of prompt compression techniques. While this provides an initial comparison, a broader investigation using our evaluation framework is needed to fully understand the trade-offs across different methods. Finally, although we examine multiple generation tasks, there is still room for further exploration, particularly in long-form QA scenarios, where prompt compression may have a more significant impact on model performance.
@@ Appendix A Implementation Details
In all the experiments we use Mistral-7B Instruct v0.2 (Jiang et al., 2023a)666huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 that contains approximately 7.24B parameters as a target LLM. We use the associated model checkpoint implemented using the PyTorch (Paszke et al., 2019) framework that is available on the HuggingFace Hub via the transformers (Wolf et al., 2020) library. We access microsoft/phi-2 (Abdin et al., 2023) [^7] https://huggingface.co/microsoft/phi-2 for LLMLingua and the Mistral 7B based xRAG (Hannibal046/xrag-7b [^8] https://huggingface.co/Hannibal046/xrag-7b) in the same way. We load the datasets from the HuggingFace using the Datasets library Lhoest et al. (2021).
@@
We train our models using two compute instances, each equipped with 8 NVIDIA A100 Tensor Core GPUs, 1152 GiB of memory, and 96 vCPUs. The training process for a new xRAG model variant typically takes around 16 hours. We adopt a batch size of 1 per GPU and leverage the Accelerate library with a gradient accumulation step of 4.999https://huggingface.co/docs/accelerate The specific hyper-parameters used during training are detailed in Table 11.
@@ Appendix D Reproducibility of Baselines
For xRAG, we assess performance on 1,000 randomly selected samples from the HotpotQA development set (see Table 12). We evaluate two versions: (i) the vanilla xRAG model released by the authors,8 and (ii) our reproduced version trained using the original data and codebase (github.com/Hannibal046/xRAG). The reproduced model follows the same pre-training and fine-tuning setup, though we use approximately 50% of the fine-tuning samples with original context paragraphs instead of retrieved ones due to limitations in reproducing the original data pre-processing steps.